OpenStudy (anonymous):

Find the limit. Use l'Hospital's Rule where appropriate. If there is a more elementary method,

OpenStudy (anonymous):

\[(\sqrt{1+3x}-\sqrt{1+7x})/x\]

hartnn (hartnn):

did you try rationalizing the numerator ?

OpenStudy (anonymous):

so taking the derivative of the top and bottom i get \[(1/(2\sqrt{1-3x}))-1/2\sqrt{1+7x}\]

OpenStudy (anonymous):

no i didnt i figure thatd be more difficult

hartnn (hartnn):

multiply numerator and denominator by \(\Large \sqrt{1+3x}+\sqrt{1+7x}\)

OpenStudy (anonymous):

ok

hartnn (hartnn):

or , once you took the derivative you can simply plug in x=0

hartnn (hartnn):

but your derivative does not seem to be correct

OpenStudy (anonymous):

one sec

OpenStudy (anonymous):

but then i still have 0 on the bottom

hartnn (hartnn):

you do ? how ??

OpenStudy (anonymous):

well it says with hospitals law you can take the derivative of the numerator and then the derivative of the top......i do. because if x approaches 0, x(√1+3x+√1+7x) equals zero

OpenStudy (aum):

You don't apply the quotient rule in L'H. You find the derivative of the numerator separately and the derivative of the denominator separately.

hartnn (hartnn):

L'Hopital's rrule say that you take the derivative of numerator and denominator separately. the denominator is x its derivative is =1

OpenStudy (anonymous):

ah i see how my derivative could be wrong

hartnn (hartnn):